OpenStudy (anonymous):

Dr. Denis drove 100km on 12L of gas. How many liters will she use to drive 25km. Show work please

OpenStudy (anonymous):

3L

OpenStudy (anonymous):

Can u see how?

OpenStudy (anonymous):

100 Km=12L 25 Km=(12/100)*25 so 3

OpenStudy (gw2011):

Set up a proportion: 100/12=25/x 100x=300 x=3

OpenStudy (anonymous):

100 km in 12L 1 km in 12/100 L 12 300 25 km in ----- * 25 L = ----- L = 3L 100 100

OpenStudy (anonymous):

yeah each liter is worth 8.25